style=”color:black”>Operations Report Feb 25th

Crew 206 Operations Report 25 – FEB – 2019

SOL: 1

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned off at 0830 (24th Feb), Turned on at
1600 (24th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 76%, Turned on Gen (24th Feb) 82%

Diesel Reading – ~ 75%
Propane Reading – ~ 25%
Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use
Water (Static tank) – around 95%, ~ 522 gallons
Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 293,7 gallons
Water (loft) – 12 marked level

Static to Loft Pump used – yes
Water Meter: 01415594 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Used
Hours: 51.5 hours
Beginning charge: 100%
Ending charge: 74%
Currently charging: Yes

Curiosity rover used: Used
Hours: 75.8 hours
Beginning charge: 100%
Ending charge: 75.5%
Currently charging: Yes

Spirit rover used: Not Used
Hours: 78.3 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: NO

Notes on rovers: Opportunity’s socket is having issues that even the reset trick doesn’t solve. Because it’s currently more viable than Spirit, I unplugged the latter to plug in Opportunity. Charge level was 76% for Opportunity in consequence.

ATV’s Used: N/A
Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used
Crew Car used and why, where? Not used

General notes and comments: We are experiencing a lot of static charges in the Hab. We don’t exactly know what is the reason behind that; but it might be a wire not properly connected to the ground.

Summary of internet: Lost Internet access twice in the day. Resetting the Nighthawk router did the trick without any issue.

Summary of suits and radios: Visor fogging on EVA#1 – check suit air supply

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Pre-built the LOAC and meteo station experiments. We ask to use one clamp to fix the LOAC for the duration of the experiment.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Propane tank still needs to be filled rapidly, how do we proceed ?

Copyright © The Mars Society. All rights reserved. | Main Site